OpenStudy (adelyn_helps):

WILL FAN AND MEDAL How much compound interest is earned on $24,500 principal, with an annual interest rate of 5% compounded annually, over 2 1/2 years?

OpenStudy (adelyn_helps):

compound interest: I = P (1 + r)t A. $30,62.50 B. $2767.82 C. $27,562.50 D. $3178.29

OpenStudy (adelyn_helps):

@AloneS

OpenStudy (adelyn_helps):

@Englishguy

OpenStudy (mathwizzard3):

First you should plug the numbers in

OpenStudy (adelyn_helps):

in where

OpenStudy (mathwizzard3):

p is 24500, r is .05, and t is 2.5

OpenStudy (mathwizzard3):

so I = 24500 (1 + .05)^2.5

OpenStudy (mathwizzard3):

Now you add in the parenthesis to get I=24500(1.05 )^2.5

OpenStudy (adelyn_helps):

I got 24,530.63

OpenStudy (adelyn_helps):

Is that correct

OpenStudy (mathwizzard3):

I got 27,678.2948877 but was t an exponent?

OpenStudy (adelyn_helps):

is that correct i mean is yours correct

OpenStudy (mathwizzard3):

Wait is t an exponent or not?

OpenStudy (adelyn_helps):

idk

OpenStudy (adelyn_helps):

yes it was actually

OpenStudy (adelyn_helps):

can you help with a few more

OpenStudy (mathwizzard3):

then I believe my answer was right :)

OpenStudy (adelyn_helps):

ok can you please help with a few more

OpenStudy (mathwizzard3):

maybe what are they?

OpenStudy (adelyn_helps):

Farris earned simple interest on $4120 with an annual interest rate of 3.4% for 9 years. How much did he earn in interest? Round your answer to the nearest cent. simple interest: I = P • r • t

OpenStudy (mathwizzard3):

okay so you would set it up like I=4120*3.4%*9

OpenStudy (mathwizzard3):

so I = 1260.72

OpenStudy (mathwizzard3):

does that make sense?

OpenStudy (adelyn_helps):

so would the answer be 1260.72

OpenStudy (mathwizzard3):

yup! :)

OpenStudy (mathwizzard3):

do you get it now?

OpenStudy (adelyn_helps):

yea

OpenStudy (mathwizzard3):

okay bye! :)

OpenStudy (adelyn_helps):

bye